Graph the following: (x+6) squared/16+ (y-7) squared/4 = 1
Answer by lwsshak3(11628)

Graph the following: (x+6) squared/16+ (y-7) squared/4 = 1
(x+6)^2/16+(y-7)^2/4=1
This is an equation of an ellipse with horizontal major axis of the standard form:
(x-h)^2/a^2+(y-k)^2/b^2=1, (a>b), (h,k)=(x,y) coordinates of center
For given equation:
center: (-6,7)
a^2=16
a=4
length of major axis=2a=8
..
b^2=4
b=2
length of minor axis=2b=4
..
see graph below:
(y-7)^2/4=1-(x+6)^2/16
(y-7)^2=4-(x+6)^2/4
(y-7)=√(4-(x+6)^2/4)
y=±(4-(x+6)^2/4)^.5+7